# Optimal Control of Nonlinear Systems with Unknown Dynamics

### v2

@@ 4.2.1 LTI System
where $[-5,-5]^{\prime}\leq\boldsymbol{x}(t)\leq[5,5]^{\prime}$ and $-1\leq u(t)\leq 1$ denote the system state and control input at time $t$, respectively.
Setup. For each simulation episode, the initial state $\boldsymbol{x}_{0}$ is uniformly sampled from the interval $[-0.1,-0.1]^{\prime}$ and $[0.1,0.1]^{\prime}$, and each episode is terminated either when $t\geq 50$. The objective of this example is to design a closed-loop optimal controller for driving the system state toward the goal state $\boldsymbol{x}_{goal}=[1,1]^{\prime}$ starting from any given initial state, for which we design the following stage cost function:
$c(\boldsymbol{x}(t),u(t))=(\boldsymbol{x}(t)-\boldsymbol{x}_{goal})^{\prime}(\boldsymbol{x}(t)-\boldsymbol{x}_{goal})+0.001u(t)^{2}.$
To conduct the simulation, we construct the deep Koopman basis function $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f}):\mathbb{R}^{2}\rightarrow\mathbb{R}^{4}$ to lift the original system states into a higher-dimensional space, which is applied in both the offline and online PGDK algorithms. We use the $ReLU$ activation function to build the $2$-layer architecture DNNs $\boldsymbol{g}$, $V$, and $\boldsymbol{\mu}$. In particular, we execute the online PGDK algorithm first to generate a data memory, which is then used as the training dataset for the offline PGDK algorithm. All DNNs are trained using the Adam optimizer [42].
Evaluation. In this simulation, we are interested in evaluating how the proposed offline and online PGDK perform on a simple control task compared to the classical optimal control technique, the linear quadratic regulator (LQR), which has access to the exact system dynamics and employs the same stage cost function. To assess the performance of the algorithms, we compare them by presenting both the simulation costs and the optimal trajectories achieved by each method, using identical initial states to ensure a fair comparison. To mitigate the influence of randomness in the DNN training, the entire experiment is repeated over $5$ independent trials.
@@
Figure 2: Learning and testing stage cost using the same initial states. Here, $\bar{c}$ denotes the averaged stage cost over each episode to account for the variance of different initial states. The solid line represents the mean stage cost across $5$ experiment trials, while the shaded region and error bars indicate the standard deviation.
@@ 4.2.2 Simulated Inverted Pendulum
where $g$, $m$, and $l$ represent the acceleration due to gravity, the mass of the pendulum, and the length of the pendulum, respectively. Here, the system state is defined as $\boldsymbol{x}(t)=[\theta(t),\dot{\theta}(t)]^{\prime}$ with the lower bound and upper bound of $[-\pi,-8]^{\prime}$ and $[\pi,8]^{\prime}$, respectively, and $u(t)$ denotes the continuous scalar torque input, constrained between $-2$ and $2$.
Setup. For each simulation episode, the initial state $\boldsymbol{x}(0)$ is sampled uniformly from the interval $[-\pi,-1]^{\prime}$ and $[\pi,1]^{\prime}$ and an episode terminates if $t>200$. The goal is to balance the pendulum to the upright position, represented by $\boldsymbol{x}_{goal}=[0,0]^{\prime}$. To achieve this, the following stage cost function is defined:
@@
To conduct the simulation, we set $\Delta t=0.02s$, and lift the system states using the deep Koopman basis function $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f}):\mathbb{R}^{2}\rightarrow\mathbb{R}^{8}$, which is utilized in both the offline and online PGDK algorithms. We use the DNNs with the same structures and activation functions as in the LTI example, except an added layer in the policy $\boldsymbol{\mu}$, which uses a $Tanh()$ activation function to enhance expressiveness for the nonlinear dynamics. The online PGDK algorithm is executed first, and the data memory generated during this process is retained and subsequently used as the training dataset for the offline PGDK implementation. We use the Adam optimizer for DNN training.
Evaluation. This simulation aims to evaluate the convergence rate of the proposed online PGDK algorithm in comparison with existing on-policy reinforcement learning (RL) methods, as well as to assess the optimality of the resulting closed-loop controller relative to model-based optimal control strategies. Following this, to assess the performance of the proposed online PGDK, we initially compare it with a representative model-free RL algorithm, the deep deterministic policy gradient (DDPG) [8], to highlight data efficiency. Subsequently, we evaluate the online PGDK alongside two additional approaches: MPC with a known exact dynamics model and MPC with a learned deep Koopman model (DKMPC) [31]. To demonstrate the optimality of the policy learned by PGDK, we compare the trial cost, $J=\sum_{t=0}^{200}c(\boldsymbol{x}(t),\boldsymbol{u}(t))$, across PGDK, MPC, and DKMPC. The comparison uses identical stage cost functions and the following seven initial states: $(\theta_{0}=\pi/12,\dot{\theta}_{0}=-1)$, $(\theta_{0}=-\pi/12,\dot{\theta}_{0}=-1)$, $(\theta_{0}=\pi/4,\dot{\theta}_{0}=1)$, $(\theta_{0}=-\pi/4,\dot{\theta}_{0}=1)$, $(\theta_{0}=\pi/2,\dot{\theta}_{0}=0)$, $(\theta_{0}=-\pi/2,\dot{\theta}_{0}=0)$, and $(\theta_{0}=\pi,\dot{\theta}_{0}=0)$. The prediction horizons for MPC and DKMPC are set to $80$ and $6$, respectively.
@@
Figure 4: Learning losses of the online (blue) and offline (black) PGDK, where $L_{f}$, $L_{J}$, and $\hat{J}_{\mu}$ represent the loss functions computed over the entire dataset at iteration $k$, while $L_{k}^{f}$, $L_{k}^{J}$, and $\hat{J}_{k}^{\mu}$ denote their corresponding values computed using the sampled data batches.
As can be seen in Fig. 5, for the inverted pendulum example, the proposed online PGDK can reach DDPG’s performance but with better data efficiency since it requires fewer episodes ($10$ episodes) to find the optimal policy compared to the DDPG ($60$ episodes).
@@
Finally, Fig. 6 shows the performance of the proposed online PGDK compared to the model-based MPC approach with the same initial states. As shown by the testing trial costs, online PGDK achieves performance close to MPC that has access to the exact system dynamics. Furthermore, online PGDK outperforms MPC when the latter relies on trajectory propagation via a learned deep Koopman model, which is susceptible to cumulative model approximation errors.
Figure 6: Testing cost of online PGDK, MPC, and DKMPC.
@@ 6 Appendix
In this section, we present the proofs of the theoretical results discussed in Section 3.3. Before delving into the main proofs, we introduce the following supporting lemma, which pertains to the convergence analysis of the TD error:
Lemma 3
[Theorem $1$, [39]] Suppose the DNN $V(\boldsymbol{x},\boldsymbol{\theta}^{J})$ is modeled using a linear function approximation as $V(\boldsymbol{x},\boldsymbol{\theta}^{J})=\boldsymbol{\phi}_{J}(\boldsymbol{x})^{\prime}\boldsymbol{\theta}^{J}$, where $\parallel\boldsymbol{\phi}_{J}(\boldsymbol{x})\parallel\leq 1$. Additionally, it is assumed that the system dynamics in (1) following the policy $\boldsymbol{\mu}$ is ergodic with a unique stationary distribution $\pi(\boldsymbol{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$. Here, for any two states $\boldsymbol{x}_{1},\boldsymbol{x}_{2}:\pi(\boldsymbol{x}_{2})=\lim_{t\rightarrow\infty}\mathbb{P}(\boldsymbol{x}_{t}=\boldsymbol{x}_{2}|\boldsymbol{x}_{0}=\boldsymbol{x}_{1})$. For $\Sigma=\sum_{\boldsymbol{x}\in\mathcal{X}}\pi(\boldsymbol{x})\boldsymbol{\phi}_{J}(\boldsymbol{x})\boldsymbol{\phi}_{J}(\boldsymbol{x})^{\prime}$, let $\omega_{J}$ be the smallest eigenvalue of $\Sigma$. Given $\omega_{J}>0$ and any constant step size $\alpha_{J}=(1-\gamma)/4$, let $\nu_{J}=(1-\gamma)^{2}\omega_{J}/4$ by following the updating rule described in (14) one has
@@ 6.1 Proof of Lemma 1
As stated in Definition 1, $\boldsymbol{\theta}^{f*}$ represents the optimal parameter satisfying ${\mathbf{L}}_{f}(A^{*},B^{*},C^{*},\boldsymbol{\theta}^{f*})\leq{\mathbf{L}}_{f}(A,B,C,\boldsymbol{\theta}^{f})$. For brevity, we define $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}=\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f})$. To establish that $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}=0$, we begin by taking the squared norm after subtracting $\boldsymbol{\theta}^{f*}$ on both sides of (12), which results in:
@@
$=$ $\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}-\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}$
$=$ $\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\!\!-2\alpha_{k}^{f}\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k},\boldsymbol{\theta}_{k}^{f}\!-\!\boldsymbol{\theta}^{f*}\rangle+\!\!\parallel\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}\!\!\!.$
The convergence proof of (22) consists of two main parts. First, we establish an upper bound for $-\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k},\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$. Then, we derive an upper bound for $\parallel\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}$.
To begin, let $\Delta\boldsymbol{\theta}_{k}^{f}=\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}$, we derive a upper bound of $-\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k},\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$ by expanding $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}$, yielding:
$\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f}),\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$ (23)
@@
$+\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}^{f*}),\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$
@@
$-\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k},\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$ (24)
$\leq$ $-\frac{1}{2N}(\parallel\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f}\parallel_{F}^{2}+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2})\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}.$
To derive an upper bound for (24), we utilize the cyclic property of the trace operator and apply the Frobenius norm bound from [43]. Specifically, for matrices $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times r}$, the following inequality holds:
$\lambda_{min}(A)Tr(B)\leq Tr(AB)\leq\lambda_{max}(A)Tr(B).$
Thus, we obtain
@@
$\geq$ $\frac{1}{2N}(\omega_{f_{1}}(Np+\!\parallel A_{k}\parallel_{F}^{2}\!+\!\parallel C_{k}\parallel_{F}^{2})\!-\!2\omega_{f_{2}}Tr(A_{k}))\eqqcolon L_{f1},$
where $\omega_{f_{1}}=\min\{\lambda_{min}(\boldsymbol{\bar{\Phi}}_{f}\boldsymbol{\bar{\Phi}}_{f}^{\prime}),\lambda_{min}(\boldsymbol{\Phi}_{f}\boldsymbol{\Phi}_{f}^{\prime})\}$ and $\omega_{f_{2}}=\lambda_{max}(\boldsymbol{\bar{\Phi}}_{f}\boldsymbol{\Phi}_{f}^{\prime})$. Here, a sufficiently large $N$ may be required to ensure $L_{f1}>0$.
To establish an upper bound for $\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}$, we proceed by utilizing the definition of ${\mathbf{L}}_{f}$, leading to the following result:
$\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}$
@@
$+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2}\parallel\bar{{\mathbf{X}}}-C_{k}\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}),$
where $[A_{k}\ B_{k}]$ and $C_{k}$ are obtained via the least squares solutions in (13). Here, given that the system states and control inputs are bounded and $\boldsymbol{g}$ is designed to be Lipschitz continuous, the norms of the matrices $A_{k}$, $B_{k}$, and $C_{k}$ from (13) remain bounded. Specifically, we define these bounds as $\parallel A_{k}\parallel\leq c_{1}$, $\parallel B_{k}\parallel\leq c_{2}$, and $\parallel C_{k}\parallel\leq c_{3}$ throughout this proof. Furthermore, we define the upper bounds of the residual vectors originating from the least squares solutions as [44] $\parallel\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-A_{k}\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-B_{k}\mathbf{U}\parallel_{F}^{2}\leq W_{1}$ and $\parallel\bar{{\mathbf{X}}}-C_{k}\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}\leq W_{2}$, which leads to following upper bound:
$\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}\leq(2(1+c_{1}^{2})W_{1}+2c_{3}^{2}W_{2})/N\coloneqq L_{f2}.$ (26)
Finally, applying the derived bounds $L_{f1}$ from (25) and $L_{f2}$ from (26) to (22), we obtain:
$\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\leq(1-2\alpha_{k}^{f}L_{f1})\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}+(\alpha_{k}^{f})^{2}L_{f2}.$ (27)
Here, to achieve $10<1-2\alpha_{k}^{f}L_{f1}<$ 1, one needs to choose step size $0<\alpha_{k}^{f}<\frac{1}{2L_{f1}}$.
Based on (27), for constant $0<\alpha_{f}<\frac{1}{2L_{f1}}$ and following the geometric series ($\sum_{k=0}^{\infty}ar^{k}=\frac{a}{1-r},|r|<1$), the following holds:
$\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}$ (28)
$\leq$ $(1-2\alpha_{f}L_{f1})^{k}\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}+\alpha_{f}^{2}L_{f2}\sum_{s=0}^{\infty}(1-2\alpha_{f}L_{f1})^{s}$
$\leq$ $(1-2\alpha_{f}L_{f1})^{k}\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}+\frac{\alpha_{f}L_{f2}}{2L_{f1}}.$
[…]
6.2 Proof of Corollary 1

### v3

@@ 4.2.1 LTI System
where $5,-5]^{\prime}\leq\boldsymbol{x}(t)\leq[5,5]^{\prime}$ and $1\leq u(t)\leq 1$ denote the system state and control input at time $t$, respectively.
Setup. For each simulation episode, the initial state $\boldsymbol{x}_{0}$ is uniformly sampled from the interval $0.1,-0.1]^{\prime}$ and $0.1,0.1]^{\prime}$, and each episode is terminated either when $t\geq 50$. The objective of this example is to design a closed-loop optimal controller for driving the system state toward the goal state $\boldsymbol{x}_{\mathrm{goal}}=[1,1]^{\prime}$ starting from any given initial state, for which we design the following stage cost function:
$c(\boldsymbol{x}(t),u(t))=(\boldsymbol{x}(t)-\boldsymbol{x}_{\mathrm{goal}})^{\prime}(\boldsymbol{x}(t)-\boldsymbol{x}_{\mathrm{goal}})+0.001u(t)^{2}.$
To conduct the simulation, we construct the deep Koopman basis function $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f}):\mathbb{R}^{2}\rightarrow\mathbb{R}^{4}$ to lift the original system states into a higher-dimensional space, which is applied in both the offline and online PGDK algorithms. To satisfy Assumption 1, the functions $\boldsymbol{g}$ and $V$ are implemented as two-layer DNNs with $ReLU$ activation functions, consisting of $400$ and $300$ neurons per layer, respectively. The policy $\boldsymbol{\mu}$ uses a $ReLU$ function with $400$ neurons in the first hidden layer and a tanhtanh function with $300$ neurons in the second hidden layer. In particular, we execute the online PGDK algorithm first to generate a data memory, which is then used as the training dataset for the offline PGDK algorithm. All DNNs are trained via the Adam optimizer [kingma2014adam].
Evaluation. In this simulation, we are interested in evaluating how the proposed offline and online PGDK perform on a simple control task compared to the classical optimal control technique, the linear quadratic regulator (LQR) with horizon length $50$, which has access to the exact system dynamics and employs the same stage cost function. For both PGDK and LQR, the system states and control inputs are clipped to satisfy state and input constraints. To assess the algorithm performance, we compare them by presenting both the simulation costs and the optimal trajectories achieved by each method, using identical initial states to ensure a fair comparison. To mitigate the influence of randomness in the DNN training, the entire experiment is repeated over $5$ independent trials.
@@
Figure 3: Learning and testing stage cost using the same initial states. Here, $\bar{c}$ denotes the averaged stage cost over each episode to account for the variance of different initial states. The solid line represents the mean stage cost across 55 experiment trials, while the shaded region and error bars indicate the standard deviation.
@@ 4.2.2 Simulated Inverted Pendulum
where $g$, $m$, and $l$ represent the acceleration due to gravity, the mass of the pendulum, and the length of the pendulum, respectively. Here, the system state is defined as $\boldsymbol{x}(t)=[\theta(t),\dot{\theta}(t)]^{\prime}$ with the lower bound and upper bound of $\pi,-8]^{\prime}$ and $\pi,8]^{\prime}$, respectively, and $u(t)$ denotes the continuous scalar torque input, constrained between 2-2 and $2$.
Setup. For each simulation episode, the initial state $\boldsymbol{x}(0)$ is sampled uniformly from the interval $\pi,-1]^{\prime}$ and $\pi,1]^{\prime}$ and an episode terminates if $t>200$. The goal is to balance the pendulum to the upright position, represented by $\boldsymbol{x}_{\mathrm{goal}}=[0,0]^{\prime}$. To achieve this, the following stage cost function is defined:
@@
To conduct the simulation, we set $\Delta t=0.02s$, and lift the system states using the deep Koopman basis function $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f}):\mathbb{R}^{2}\rightarrow\mathbb{R}^{8}$, which is utilized in both the offline and online PGDK algorithms. The functions $\boldsymbol{g}$ and $V$ adopt the same architectures and activation functions as in the LTI example. The policy $\boldsymbol{\mu}$ is implemented as a two-layer DNN with $ReLU$ functions in the hidden layers ($400$ and $300$ neurons, respectively) and a tanhtanh output layer. The online PGDK algorithm is executed first, and the data memory generated during this process is retained and subsequently used as the training dataset for the offline PGDK implementation. We use the Adam optimizer for DNN training.
Evaluation. This simulation evaluates the convergence rate of the proposed online PGDK algorithm (trained for $200$ episodes) compared with existing on-policy RL methods and assesses the optimality of the resulting closed-loop controller relative to model-based optimal control strategies. We compare PGDK with four additional approaches:
  • DDPG: A representative actor-critic RL algorithm trained for $1000$ episodes using the same actor–critic learning rates and optimizer as PGDK.
  • MPC: Model predictive control with the exact system dynamics, horizon length $50$, and input constraints applied during optimization.
  • DKMPC-v1: MPC using a learned deep Koopman model trained offline using $9000$ state–input pairs generated by applying control inputs sampled from a normal distribution [dk], with horizon $8$ and input constraints.
  • DKMPC-v2: MPC using the dynamics learned by PGDK, horizon $20$, with input constraints.
For PGDK and DDPG, states and inputs are clipped to satisfy constraints. Policy optimality is evaluated via the trial cost $J=\sum_{t=0}^{200}c(\boldsymbol{x}(t),\boldsymbol{u}(t))$, across all methods. The comparison uses identical stage cost functions and seven initial states, ordered from closest to farthest from the goal: $\theta_{0},\dot{\theta}_{0}]=[\pi/12,-1]$, $\pi/12,-1]$, $\pi/4,1]$, $\pi/4,1]$, $\pi/2,0]$, $\pi/2,0]$, $\pi,0]$.
@@
Figure 5: Learning losses of the online (blue) and offline (black) PGDK, where $L_{f}$, $L_{J}$, and $\hat{J}$ represent the loss functions computed over the entire dataset at iteration $k$, while $L_{k}^{f}$, $L_{k}^{J}$, and $\hat{J}_{k}^{\mu}$ denote their corresponding values computed using the sampled data batches.
As can be seen in Figs. 6-7, for the inverted pendulum example, the proposed online PGDK can reach DDPG’s performance but with better data efficiency since it requires fewer episodes ($10$ episodes) to find the optimal policy compared to the DDPG ($60$ episodes).
@@
Fig. 7 further compares online PGDK with model-based MPC under identical initial states. The testing trial costs show that both online PGDK and DDPG approach the performance of MPC with access to exact system dynamics. Notably, online PGDK outperforms MPC when the latter relies on trajectory propagation via deep Koopman models (both DKMPC-v1 and DKMPC-v2), which are prone to cumulative model approximation errors.
Figure 7: Testing cost plotted on a logarithmic scale.
@@ 6 Appendix
In this section, we present the proofs of the theoretical results discussed in Section 3.3.
@@ 6.1 Proof of Lemma 1
For brevity, we denote $\Delta\boldsymbol{\theta}_{k}^{f}=\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}$ and $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}=\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f})$. To establish that $\lim_{k\rightarrow\infty}\parallel\Delta\boldsymbol{\theta}_{k}^{f}\parallel^{2}=0$, we begin by taking the squared norm after subtracting $\boldsymbol{\theta}^{f*}$ on both sides of (12), which results in:
@@
$=$ $\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}-\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}\parallel^{2}$
$=$ $\parallel\Delta\boldsymbol{\theta}_{k}^{f}\parallel^{2}\!\!-2\alpha_{k}^{f}\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f},\Delta\boldsymbol{\theta}_{k}^{f}\rangle+\!\!\parallel\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}\parallel^{2}\!\!\!.$
The convergence proof of (21) consists of two main parts. First, we establish an upper bound for $\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f},\Delta\boldsymbol{\theta}_{k}^{f}\rangle$. Then, we derive an upper bound for $\parallel\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}\parallel^{2}$.
To this end, we derive a upper bound of $\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f},\Delta\boldsymbol{\theta}_{k}^{f}\rangle$ by expanding $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}$ and using $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A^{*},B^{*},C^{*},\boldsymbol{\theta}^{f*})=0$, yielding:
$\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f}),\Delta\boldsymbol{\theta}_{k}^{f}\rangle$ (22)
@@
$+\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}^{f*}),\Delta\boldsymbol{\theta}_{k}^{f}\rangle$
@@
$-\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f},\Delta\boldsymbol{\theta}_{k}^{f}\rangle$ (23)
$\leq$ $-\frac{1}{2N}(\parallel\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f}\parallel_{F}^{2}+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2})\parallel\Delta\boldsymbol{\theta}_{k}^{f}\parallel^{2}.$
To derive an upper bound for (23), we utilize the cyclic property of the trace operator and apply the Frobenius norm bound and the matrix Cauchy–Schwarz inequality, which yields
@@
$\geq$ $\frac{1}{2N}(\omega_{f}(Np+\!\parallel A_{k}\parallel_{F}^{2}\!+\!\parallel C_{k}\parallel_{F}^{2})\!-\!2\sqrt{N}\parallel A_{k}\parallel_{F})\eqqcolon L_{f1},$
where $\omega_{f}=\min\{\lambda_{\mathrm{min}}(\boldsymbol{\bar{\Phi}}_{f}\boldsymbol{\bar{\Phi}}_{f}^{\prime}),\lambda_{\mathrm{min}}(\boldsymbol{\Phi}_{f}\boldsymbol{\Phi}_{f}^{\prime})\}$.
To establish an upper bound for $\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}\parallel^{2}$, we proceed by utilizing the definition of ${\mathbf{L}}_{f}$, leading to the following result:
$\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}\parallel^{2}$
@@
$+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2}\parallel\bar{{\mathbf{X}}}-C_{k}\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}).$
Given that $A_{k}\ B_{k}]$ and $C_{k}$ are obtained via the least squares solutions in (13), the system states and control inputs are bounded, and $\boldsymbol{g}$ is Lipschitz continuous, the norms of the matrices $A_{k}$, $B_{k}$, and $C_{k}$ (13) remain bounded. We denote these bounds as: $\parallel A_{k}\parallel\leq c_{1}$, $\parallel B_{k}\parallel\leq c_{2}$, and $\parallel C_{k}\parallel\leq c_{3}$ throughout this proof. Furthermore, let the residuals from the least squares solutions be bounded as [van1991total]: $\parallel\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-A_{k}\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-B_{k}\mathbf{U}\parallel_{F}^{2}\leq W_{1}$, $\parallel\bar{{\mathbf{X}}}-C_{k}\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}\leq W_{2}$, which leads to following upper bound:
$\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{k}^{f}\parallel^{2}\leq(2(1+c_{1}^{2})W_{1}+2c_{3}^{2}W_{2})/N\coloneqq L_{f2}.$ (25)
Finally, applying the derived bounds $L_{f1}$ from (24) and $L_{f2}$ from (25) to (21), and using $1-x\leq e^{-x}$ for real $x$, yields:
$\parallel\Delta\boldsymbol{\theta}_{k+1}^{f}\parallel^{2}$ (26)
$\leq$ $(1-2\alpha_{k}^{f}L_{f1})\parallel\Delta\boldsymbol{\theta}_{k}^{f}\parallel^{2}+(\alpha_{k}^{f})^{2}L_{f2}$
$\leq$ $\prod_{i=0}^{k}(1-2\alpha_{i}^{f}L_{f1})\parallel\Delta\boldsymbol{\theta}_{0}^{f}\parallel^{2}+L_{f2}\sum_{j=0}^{k}(\alpha_{j}^{f})^{2}\!\!\!\prod_{i=j+1}^{k}(1-2\alpha_{i}^{f}L_{f1})$
$\leq$ $\parallel\Delta\boldsymbol{\theta}_{0}^{f}\parallel^{2}e^{-2L_{f1}\sum_{i=0}^{k}\alpha_{i}^{f}}+L_{f2}\sum_{j=0}^{k}(\alpha_{j}^{f})^{2}\!\!\!\prod_{i=j+1}^{k}(1-2\alpha_{i}^{f}L_{f1}).$
Hence, to achieve $\lim_{k\rightarrow\infty}\parallel\Delta\boldsymbol{\theta}_{k}^{f}\parallel^{2}=0$, one needs to choose step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}^{f}=\infty$ and $\sum_{k=0}^{\infty}(\alpha_{k}^{f})^{2}<\infty$.
@@
Moreover, applying constant $0<\alpha_{f}<\frac{1}{2L_{f1}}$ to (26) and using the geometric series formula ($\sum_{k=0}^{\infty}ar^{k}=\frac{a}{1-r},|r|<1$), we obtain:
$\parallel\Delta\boldsymbol{\theta}_{k}^{f}\parallel^{2}$ (29)
$\leq$ $(1-2\alpha_{f}L_{f1})^{k}\parallel\Delta\boldsymbol{\theta}_{0}^{f}\parallel^{2}+\alpha_{f}^{2}L_{f2}\sum_{s=0}^{\infty}(1-2\alpha_{f}L_{f1})^{s}$
$\leq$ $(1-2\alpha_{f}L_{f1})^{k}\parallel\Delta\boldsymbol{\theta}_{0}^{f}\parallel^{2}+\frac{\alpha_{f}L_{f2}}{2L_{f1}}.$